OpenStudy (anonymous):

a boy p stands on top of a tower another boy q throws a ball vertically up from the ground so that it just reaches p the time taken is t0 the speed with which p should throw the ball vertically downwards so that it reaches the ground in time t0/2??

OpenStudy (anonymous):

Do we assume that the boy who throws the boy has some height or that the ball just comes off the ground?

OpenStudy (dumbcow):

i get an answer of v = 24*t_0 @santistebanc , can you verify that for me

OpenStudy (anonymous):

How did you get that answer?

OpenStudy (dumbcow):

using the height function \[h(t) = -16t^{2}+v_0t + h_0 sorry equation editor not working for me

OpenStudy (anonymous):

hey the options are (a) gt0 (b) 2gt0 (c) 3/4gt0 (d) gt0/4

OpenStudy (anonymous):

where g is gravity/acceleration right?

OpenStudy (dumbcow):

oh they leave it in terms of g 24 = (3/4)*32 --> 3/4g*t_0

OpenStudy (anonymous):

yeah that's what I got too. I just use one of the kinematic equaions. The one where you only vi and no vf.

OpenStudy (anonymous):

@ dumbcow you are right, the correct answer is C

OpenStudy (anonymous):

i dont think I understood the problem

OpenStudy (anonymous):

Thanks!)

OpenStudy (anonymous):

can anyone give a detailed equation

OpenStudy (anonymous):

i cant understand the one which is written i know that answer is correct

OpenStudy (anonymous):

Let's wait for dumbcow to answer.

OpenStudy (anonymous):

dont you have to consider the difference of height between the two boys?

OpenStudy (anonymous):

no

OpenStudy (dumbcow):

the height function h(t) = -g/2 t^2 +v_0t +h_0 q throws it up to p such that it reaches highest point with initial velocity v_0 in time t h_0 = 0 highest point is when velocity = 0 v(t) = -gt +v_0 = 0 --> v_0 = gt final height is H --> H = -g/2 t^2 +gt^2 --> H = g/2 t^2 Now set up 2nd throw, which has negative initial velocity and takes time t/2 to reach same height H initial height is now H final height is 0 0 = -g/2 (t/2)^2 -v(t/2) +H substitute in for H 0 = -g/2 (t/2)^2 -v(t/2) + g/2 t^2 solve for v

OpenStudy (anonymous):

can some one use kinematics equations to solve this

OpenStudy (dumbcow):

is the height function not a kinematic equation? sorry not a physics guy

OpenStudy (anonymous):



OpenStudy (anonymous):

I assume that all of you know the standard defintions and laws and the variables used Basically we can divide this problem in 2 parts Part 1- ball from ground to tower Let "s" be the height of the tower First using this equation \[v = u + at\] I get \[0 = u - g t_0\] \[t_0 = u/g\] Also \[s = ut + \frac{1}{2} at^2\] \[s =\frac{u^2 }{g} -\frac{1}{2} g \frac{u^2}{g^2}\] \[s =\frac{u^2 }{g} -\frac{1}{2} \frac{u^2}{g}\] \[s =\frac{1}{2} \frac{u^2}{g}\] Part 2--> Tower to ground We know \[s =\frac{1}{2} \frac{u^2}{g}\] Again using the following equation \[s = ut + \frac{1}{2} at^2\] We get \[\frac{1}{2} \frac{u^2}{g} = u't' + \frac{1}{2}g t'^2\] Now we know that\[ t ' = \frac{t_0}{2}= \frac{u}{2g}\] Therefore \[\frac{1}{2} \frac{u^2}{g} = u'\frac{u}{2g}+ \frac{1}{2}g \frac{u^2}{(2g)^2}\] \[\frac{1}{2} \frac{u^2}{g} = u'\frac{u}{2g}+ \frac{1}{8} \frac{u^2}{g}\] \[\frac{3}{8} \frac{u^2}{g} = u'\frac{u}{2g}\] \[\frac{3}{8} \frac{u}{g} = u'\frac{1}{2g}\] \[\frac{3}{4} {u} = u'\] Therefore the boy P should throw the ball down with the velocity which is equal to 3/4 times with which the boy Q threw the ball up

OpenStudy (anonymous):

I am sorry if I have given the full answer but it is to dispense all doubts related to the solution. I have reviewed the solution on my side but if you find any mistakes , do notify me

OpenStudy (anonymous):

thanxxx

OpenStudy (anonymous):

Welcome. I forgot the last step .. Substitute u = gt_0 and you will get your answer